OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

first divide by -4 so abs( x + 5) = 4 so x + 5 = 4 and x+ 5 = -4 solve for x:)

OpenStudy (anonymous):

x = −32 and x = −4?

OpenStudy (anonymous):

what how did you get - 32

OpenStudy (anonymous):

there are 4 answer choices ill post

OpenStudy (anonymous):

x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

i know its not C @tejasvir

OpenStudy (anonymous):

all you have to do is solve for x in both of these equations x+ 5 = 4 and x+5 = -4 you should get -1 and -9 as your answers

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yea anytime:)